# Animate Your Thoughts: Decoupled Reconstruction of Dynamic Natural Vision from Slow Brain Activity

### v1

@@ 2.4 Feature-to-video Stage
Inflated Stable Diffusion for Video Reconstruction
Despite the rapid development of video generation models capable of producing vivid videos from text conditions, it is crucial to emphasize that the objective of our project is to disentangle semantic, structural, and motion information from fMRI to fully reconstruct the stimulus video. Utilizing pre-trained video generation models could obscure whether the motion information in the reconstructed video originates from the fMRI or external video data.
To address this issue, we employ the network inflation[46, 47, 39] technique to implement an inflated Stable Diffusion, which is used to reconstruct each frame of the video without introducing additional motion information. For further details, please refer to the Appendix B.2.
@@ 3.1 Datasets
In this study, we utilize three publicly available video-fMRI datasets, which encompass paired stimulus videos and their corresponding fMRI responses. As depicted in Table 2, these datasets collectively comprise brain signals recorded from multiple healthy subjects while they are viewing the videos. The video stimuli are diverse, covering animals, humans, and natural scenery. For detailed information on the datasets and preprocessing steps, please refer to Appendix C.
Dataset Adopted participants TR Train samples Test samples
CC2017[31] 3 2s 4320 1200
HCP[48] 3 1s 2736 304
Algonauts2021[49] 10 1.75s 900 100
Table 2: Characteristics of the video-fMRI datasets used in our experiments
@@ 3.2 Evaluation Metrics
To comprehensively and fairly evaluate the performance of our model, we propose the following evaluation metrics.
Semantic-level metrics
Following prior studies[19, 36], we use the N-way top-K accuracy classification test and VIFI-score as the semantics-level metrics. For the classification test, we compare the ground truth (GT) against the predicted video (PV) classifications using a classifier. A trial is successful if the GT class ranks within the top-K probabilities from the PV’s classification among N randomly selected classes. Additionally, we implement two modes: image-based (2-way-I) and video-based (2-way-V). We describe this evaluation method in Algorithm 2. For the VIFI-score, we utilize VIFICLIP[53]—a CLIP model fine-tuned on the video dataset—to extract features from both the GT and the PV, followed by the calculation of cosine similarity.
Pixel-level metrics
We employ the structural similarity index measure (SSIM), peak signal-to-noise ratio (PSNR), and hue-based Pearson correlation coefficient (Hue-pcc) as pixel-level metrics.
Spaciotemporal (ST) -level metric
We adopt the CLIP-pcc, a common metric in the field of video editing, which involves computing CLIP image embeddings on each frame of the predicted videos and reporting the average cosine similarity between all pairs of adjacent video frames.
@@ 4 Results
Figure 4: Reconstruction results of CC2017 dataset. Our reconstructed results are highlighted with red boxes, while those of Wen and Nishimoto are delineated by blue and green boxes, respectively.
@@ 4.1 Comparative Experimental Results
We compare our model with all previous video reconstruction models [^1] It should be noted that when replicating the results of Nishimoto et al.[28] on the CC2017 dataset, we utilized videos from the training sets of both CC2017 and HCP as the natural movie prior. on the CC2017 dataset. Visual comparisons are presented in Figure 4, while quantitative comparisons are detailed in Table 3. In the computation of quantitative metrics, the results of Wen et al.[31] pertain to the first segment of the test set, whereas the results of other researchers are derived from the whole test set. The findings on HCP and Algonauts2021 datasets are elaborated in Appendices E.2 and E.3, respectively.
Semantic-level $\uparrow$ Pixel-level $\uparrow$ ST-level $\uparrow$
2-way-I 2-way-V VIFI-score SSIM PSNR Hue-pcc CLIP-pcc
Nishimoto[28] 0.727±$\pm 0.04$ —— —— 0.116±$\pm 0.09$ 8.012±$\pm 2.31$ 0.753±$\pm 0.12$ ——
Wen[31] 0.758±$\pm 0.03$ —— —— 0.114±$\pm 0.15$ 7.646±$\pm 3.48$ 0.647±$\pm 0.11$ ——
Wang[32] 0.713±$\pm 0.04$ 0.773±$\pm 0.03$ 0.596±$\pm 0.07$ 0.118±$\pm 0.08$ 11.432±$\pm 2.42$ 0.589±$\pm 0.18$ 0.402±$\pm 0.41$
Kupershmidt[34] 0.764±$\pm 0.03$ 0.771±$\pm 0.03$ 0.585±$\pm 0.08$ 0.135±$\pm 0.08$ 8.761±$\pm 2.22$ 0.606±$\pm 0.14$ 0.386±$\pm 0.47$
Chen[36] 0.792±$\pm 0.03$ 0.853±$\pm 0.03$ 0.587±$\pm 0.08$ 0.171±$\pm 0.08$ 8.662±$\pm 1.52$ 0.760±$\pm 0.10$ 0.408±$\pm 0.46$
Ours(sub1) 0.809±$\pm 0.03$ 0.837±$\pm 0.02$ 0.602±$\pm 0.07$ 0.301±$\pm 0.09$ 9.134±$\pm 1.48$ 0.768±$\pm 0.12$ $\pm 0.42$
Ours(sub2) 0.804±$\pm 0.29$ 0.832±$\pm 0.03$ 0.604±$\pm 0.08$ 0.287±$\pm 0.11$ 9.049±$\pm 1.45$ 0.795±$\pm 0.12$ $\pm 0.42$
Ours(sub3) 0.792±$\pm 0.03$ 0.833±$\pm 0.03$ 0.600±$\pm 0.08$ 0.349±$\pm 0.11$ 9.306±$\pm 1.54$ 0.791±$\pm 0.12$ 0.415±$\pm 0.39$
Table 3: Quantitative comparison of reconstruction results. All metrics indicate superior performance with larger values, with the best results highlighted in bold and the second-best results underlined.
Table 3 indicates that our model achieves SOTA performance in 5 out of 7 metrics, securing the second place in the remaining two. Specifically, our model outperforms the previous SOTA model by 76% in terms of SSIM, which underscores the benefits of incorporating structural information. Specifically, as depicted in Figure 4, our reconstruction results contain richer semantic information compared to earlier models, such as a girl and a yellow dog being held in someone’s arms. In contrast to Mind-video by Chen et al.[36], our results are more consistent with the ground truth in terms of fine-grained structural and motion information. For instance, the reconstructed girl’s clothing color, the dog’s fur color, and the positioning of the forest along the coastline are closer to the stimulus videos. Regarding motion information, our results depict the dog being petted and a noticeable camera movement in the coral reef scene. Additional reconstruction results on other subjects, as well as instances of reconstruction failure, are presented in Appendix E.1.
@@ 4.2 Ablation Study
Semantic-$\uparrow$ Pixel-$\uparrow$ ST-level $\uparrow$
2-way-I 2-way-V VIFI-score SSIM PSNR Hue-pcc CLIP-pcc
w/o Semantic 0.679±$\pm 0.04$ 0.766±$\pm 0.04$ 0.523±$\pm 0.07$ $\pm 0.09$ 8.005±$\pm 1.57$ 0.737±$\pm 0.11$ 0.123±$\pm 0.31$
w/o Structure 0.789±$\pm 0.03$ 0.814±$\pm 0.03$ 0.555±$\pm 0.08$ 0.184±$\pm 0.08$ 8.712±$\pm 1.37$ 0.791±$\pm 0.11$ 0.260±$\pm 0.41$
w/o Motion 0.674±$\pm 0.04$ 0.789±$\pm 0.03$ 0.585±$\pm 0.08$ $\pm 0.13$ 8.611±$\pm 2.43$ 0.715±$\pm 0.14$ 0.376±$\pm 0.42$
Full Model 0.809±$\pm 0.03$ 0.837±$\pm 0.02$ 0.602±$\pm 0.07$ 0.301±$\pm 0.10$ 9.134±$\pm 1.51$ 0.768±$\pm 0.11$ 0.425±$\pm 0.41$
Table 4: Ablation study on our proposed decoders. 100 repetitions are conducted on the metrics 2-way-I and 2-way-V, while 5 trials are performed on other metrics, with the results being averaged across all samples in test set and trials. Colors reflect statistical significance (paired t-test) compared to the Full Model. $p<0.0001$ (purple); $p<0.01$ (pink); $p<0.05$ (yellow); $p>0.05$ (green).
In this subsection, we conduct a detailed ablation study to assess the effectiveness of the three decoders we proposed and to evaluate the impact of various hyperparameters on video reconstruction (See Appendix E.4).
First, we present the results obtained using the full model. Then, on the basis of the full model, we separately eliminate the semantic decoder (w/o Semantic) and the structure decoder (w/o Structure) by replacing their outputs with random noise. For the consistency motion generator, we replaced it with 8 simple MLPs to model each frame individually (w/o Motion). Table 4 demonstrates that the removal of any decoder results in a significant decline in the model’s performance across nearly all metrics, which shows the efficacy of our proposed decoders.
@@ 5.1 Have we truly decoded motion information from fMRI?
This work focuses on the video reconstruction from fMRI, aiming for motion consistency between the reconstructed and stimulus videos. We specifically design a Consistency Motion Generator (CMG) to decode motion information from fMRI. Following the work of Wang et al. [32], we perform a permutation test on 3 subjects from the CC2017 dataset to ascertain whether this module decodes the correct motion information from fMRI. Specifically, for each 8-frame reconstructed video clip from each subject, we randomly shuffle the frame order 100 times and compute pixel-level and spaciotemporal-level metrics between the actual and shuffled frames. Subsequently, we estimate the P-value by the following formula: $P=\sum_{i=1}^{100}\delta_{i}/100$, where $\delta_{i}=1$ if the $i$th permutation outperforms the reconstruction result in the original order based on the metrics; otherwise, $\delta_{i}=0$. A lower P-value signifies a closer alignment between the sequential order of the reconstructed video and the ground truth. We repeat the permutation test 5 times under conditions with and without the CMG, as illustrated in Figure 5. It can be observed that the P-value significantly increased across nearly all metrics for all subjects when the CMG is removed, suggesting that we truly decodes motion information from fMRI.
Figure 5: The result of permutation test on the CC2017 dataset. The experiment is repeated 5 times on 3 subjects, with the mean and std presented in subplots (a), (b), and (c), respectively. Paired t-tests are performed, with significance denoted as $p<0.001(***)$, $p<0.01(**)$, $p<0.05(*)$, and $p>0.05(NS)$ for non-significant results.
@@ 5.2 Which brain regions are responsible for decoding different features, respectively?
To investigate voxels in which brain regions are responsible for decoding different features (semantic, structure, motion) during the fMRI-to-feature stage, we compute the voxel-wise importance maps in the visual cortex. Specifically, for a trained decoder, we multiply the weight matrix of the linear layers, then average the result across the feature dimension, and normalize it to estimate the importance weight for each voxel. A higher weight indicates that the voxel plays a more significant role in feature decoding. We project the importance maps of subject 1’s voxels from the CC2017 dataset onto the visual cortex, as depicted in Figure 6. To obtain ROI-wise importance maps, we calculate the average of the importance weights of voxels contained within each Region of Interest (ROI), with the results presented in Figure 7. The results from other subjects are presented in Appendix E.5.
@@
Figure 6 (a) indicates that high-level visual cortex (HVC, such as MT, MST and TPOJ) contribute more significantly to the decoding of semantic feature, with a calculated weight of 2.588, accounting for 60.5% of the total, as shown in Figure 7 (a). In contrast, low-level visual cortex (LVC, such as V1, V2, V3) have a weight of 1.685, representing 39.5%. Although it is not immediately apparent from Figure 6 (b) which ROI contributes most to the decoding of structural feature, Figure 7 (b) reveals that V1 and V2 have the greatest weight, with HVC having a weight of 1.279 (36.03%), and LVC weighing 2.271 (63.97%). Considering the aforementioned findings, our results are plausible from the perspective of cognitive neuroscience. It is generally believed that LVC is predominantly responsible for processing low-level information of visual stimuli [4, 54, 55], such as orientation and contour. Meanwhile, V4 is involved in color processing [56, 57]. In contrast, HVC is responsible for processing high-level semantic information of objects[58], including category.
@@
Figure 6 (c) indicates that both LVC and HVC contribute to the decoding of motion information, with significant weight attributed to V1 and MT. As derived from Figure 7 (c), the weight distribution between LVC and HVC is comparable, accounting for 42.4% and 57.6%, respectively. This observation is consistent with previous work[59], which validates the function of MT in visual motion processing. Furthermore, our findings affirm that the spatial and temporal modules designed in CMG effectively capture spatiotemporal information from across both LVC and HVC.

### v2

@@ 3.3 Feature-to-video Stage
Inflated Stable Diffusion for Video Reconstruction.
Despite the rapid development of video generation models capable of producing vivid videos from text conditions, it is crucial to emphasize that the objective of our work is to disentangle semantic, structural, and motion information from fMRI to reconstruct the stimulus video. Utilizing pre-trained video generation models could obscure whether the motion information in the reconstructed video originates from the fMRI or external video data.
To address this issue, we employ the network inflation (Carreira & Zisserman (2017); Khachatryan et al. (2023); Wu et al. (2023)) technique to implement an inflated Stable Diffusion, which is used to reconstruct each frame of the video without introducing additional motion information. Specifically, after the motion features $\Phi(\mathbf{v}_{i})\in\mathbb{R}^{B\times f\times 3\times\frac{H}{8}\times\frac{W}{8}}$ are decoded, they are reshaped ($(B,f,3,\frac{H}{8},\frac{W}{8})\rightarrow(B\cdot f,3,\frac{H}{8},\frac{W}{8})$) and input into the U-Net of Stable Diffusion for reverse denoising. The result is then mapped back to pixel space through the VQ-VAE decoder and reshaped ($(B\cdot f,3,H,W)\rightarrow(B,f,3,H,W)$) to yield the final video $\mathbf{v}_{i}\in\mathbb{R}^{B\times f\times 3\times H\times W}$. In this context, $B$ denotes the batch dimension, with $B=1$ during inference.
@@ 4.1 Datasets
In this study, we utilize three publicly available video-fMRI datasets, which encompass paired stimulus videos and their corresponding fMRI responses. As depicted in Table 1, these datasets collectively comprise brain signals recorded from multiple healthy subjects while they are viewing the videos. The video stimuli are diverse, covering animals, humans, and natural scenery. For detailed information on the datasets and preprocessing steps, please refer to Appendix B.
Table 1: Characteristics of the video-fMRI datasets used in our experiments.
Dataset Adopted subjects TR Train samples Test samples
CC2017 (Wen et al. (2018)) 3 2s 4320 1200
HCP (Marcus et al. (2011)) 3 1s 2736 304
Algonauts2021 (Cichy et al. (2021)) 10 1.75s 900 100
@@ 4.2 Evaluation Metrics
To comprehensively evaluate the performance of our model, we use the following evaluation metrics.
Semantic-level metrics.
Following prior studies (Chen et al. (2023; 2024)), we use the N-way top-K accuracy classification test and VIFI-score as the semantic-level metrics. For the classification test, we implement two modes: image-based (2-way-I) and video-based (2-way-V). We describe this evaluation method in Algorithm 2. For the VIFI-score, we utilize a CLIP model fine-tuned on the video dataset (VIFICLIP) (Rasheed et al. (2023)) to extract features from both the ground truth and predicted videos, followed by the calculation of cosine similarity.
Pixel-level metrics.
We employ the structural similarity index measure (SSIM), peak signal-to-noise ratio (PSNR), and hue-based Pearson correlation coefficient (Swain & Ballard (1991)) (Hue-pcc) as pixel-level metrics.
Spatiotemporal (ST) -level metrics.
We adopt CLIP-pcc, a widely used metric in video editing research (Wu et al. (2023)), to evaluate the smoothness and consistency between consecutive video frames. This metric computes the CLIP image embeddings for each frame in the predicted videos and reports the average cosine similarity between all pairs of adjacent frames. Considering the input to the video reconstruction task contains substantial noise, there are instances where every pixel of each reconstructed video frame is either zero or noise, which would artificially inflate the CLIP score if used directly. Therefore, we calculate the CLIP score only when the VIFI-CLIP value exceeds the average level (0.6); otherwise, we assign a score of 0.
To measure the similarity of motion trajectories, we introduce End-Point Error (EPE) (Barron et al. (1994)), which calculates the Euclidean distance between the predicted and ground truth endpoints for each frame.
@@ 5.1 Comparative Experimental Results
Figure 5: Reconstruction results on CC2017 dataset. Our reconstructed results are highlighted with a red box, while those of Wen and Nishimoto are delineated by blue and green boxes, respectively.
Models Training data Semantic-level ↑ Pixel-level ↑ ST-level
2-way-I 2-way-V VIFI SSIM PSNR Hue-pcc CLIP↑ EPE↓
Nishimoto (Nishimoto et al. (2011)) CC2017 0.742 —— —— 0.119 8.383 0.737 —— ——
Wen (Wen et al. (2018)) CC2017 0.771 —— —— 0.130 8.031 0.637 —— ——
Kupershmidt (Kupershmidt et al. (2022)) CC2017+EV 0.769 0.768 0.591 0.140 10.637 0.616 0.382 ——
f-CVGAN (Wang et al. (2022)) CC2017 0.721 0.777 0.592 0.108 11.043 0.583 0.399 6.344
Mind-video $\dagger$ (Chen et al. (2024)) CC2017+HCP 0.797 0.848 0.593 0.177 8.868 0.768 0.409 6.125
Ours CC2017 0.805 0.830 0.608 0.321 9.220 0.786 0.425 5.422
Table 2: Quantitative comparison of reconstruction results on the CC2017 dataset. All metrics are averaged across all samples and three subjects, with the best results highlighted in bold and the second-best results underlined. EV refers to the External Video Dataset. The symbol $\dagger$ refers to using Stable Diffusion fine-tuned on video data. 100 sample sets were constructed using the bootstrap method for hypothesis testing. Colors reflect statistical significance (Wilcoxon test for paired samples) compared to our model. $p<0.0001$ (purple); $p<0.01$ (pink); $p<0.05$ (yellow); $p>0.05$ (green).
@@
We compare our model with all previous video reconstruction models on the aforementioned datasets. In the computation of quantitative metrics, the results of Wen et al. (2018) pertain to the first segment of the test set, whereas the results of other researchers are derived from the whole test set. Visual comparisons on CC2017 dataset are presented in Figure 5, while quantitative comparisons are detailed in Table 2, which indicates that our model achieves SOTA performance in six out of eight metrics. Specifically, our model outperforms the previous SOTA model by 83% and 13% in terms of SSIM and EPE respectively, which underscores the benefits of incorporating structural and motion information. The results in Tables 4 and 4 demonstrate that our model maintains strong performance on other datasets as well. For instance, our model outperforms Mind-video by 196%/ 21%/ 4% on the HCP and 275%/ 27%/ 5% on the Algonauts 2021 dataset across the three pixel-level metrics.
[…]
Models Semantic-level ↑ Pixel-level ↑
2-way-I SSIM PSNR Hue-pcc
Nishimoto 0.658 0.321 11.316 0.645
Wen 0.702 0.058 10.197 0.727
f-CVGAN —— 0.159 13.033 ——
Mind-video 0.779 0.116 9.275 0.793
Ours 0.786 0.344 11.233 0.829
Table 3: Quantitative comparison of reconstruction results on the HCP dataset. The full table can be found in Appendix E.7.
[…]
Models Semantic-level ↑ Pixel-level ↑
2-way-I SSIM PSNR Hue-pcc
Nishimoto 0.687 0.443 9.578 0.666
Wen 0.625 0.172 8.822 0.627
Mind-video 0.681 0.124 8.673 0.796
Ours 0.701 0.465 10.989 0.833
Table 4: Quantitative comparison of reconstruction results on the Algonauts 2021 dataset. The full table can be found in Appendix E.8.
Dataset CC2017
Subjet1 Subjet2 Subjet3 Average
Model Test set top-10 top-100 top-10 top-100 top-10 top-100 top-10 top-100
Wen (Wen et al. (2018)) Small $2.17_{*}$ $19.50_{*}$ $3.33_{*}$ $19.17_{*}$ —— —— $2.75_{*}$ $19.33_{*}$
Kupershmidt (Kupershmidt et al. (2022)) Small $1.09_{*}$ $8.57_{*}$ $0.92_{*}$ $8.24_{*}$ $0.84_{*}$ $8.24_{*}$ $0.95_{*}$ $8.35_{*}$
Mind-video (Chen et al. (2024)) Small $\textbf{3.22}_{*}$ $19.08_{*}$ $2.75_{*}$ $16.83_{*}$ $3.58_{*}$ $22.08_{*}$ $3.18_{*}$ $19.33_{*}$
[…]
Wen (Wen et al. (2018)) Large $1.41_{*}$ $11.58_{*}$ $2.08_{*}$ $9.58_{*}$ —— —— $1.75_{*}$ $10.58_{*}$
Kupershmidt (Kupershmidt et al. (2022)) Large $0.17_{*}$ $2.94_{*}$ $0.17_{*}$ $2.77_{*}$ $0.25_{*}$ $2.18_{*}$ $0.19_{*}$ $2.63_{*}$
Mind-video (Chen et al. (2024)) Large $1.75_{*}$ $7.17_{*}$ $0.83_{*}$ $5.17_{*}$ $1.25_{*}$ $9.00_{*}$ $1.28_{*}$ $7.11_{*}$
[…]
Table 5: Retrieval accuracy (%) on CC2017 dataset. For the ’small test set’, the chance-level accuracies for top-10 and top-100 accuracy are 0.83% and 8.3%, respectively. For the ’large test set’, the chance-level accuracies for top-10 and top-100 accuracy are 0.24% and 2.4%, respectively. Five random seeds were used during the training of the semantic decoder, and the results were tested for statistical significance. * denotes our performance is significantly better than the compared method (Wilcoxon test, p<0.05).
In addition to the reconstruction task, we evaluate the retrieval task on the CC2017 dataset. We use top-10 accuracy and top-100 accuracy as evaluation metrics. To assess the model’s generalization capability, we perform retrieval not only on the CC2017 test set with 1,200 samples (’Small’) but also on an extended stimulus set. Specifically, we augment the collection with 3,040 video clips from the HCP dataset, resulting in a total of 4,240 samples (’Large’).
As shown in Table 5, our model achieves superior performance across all three subjects in the CC2017 dataset. Compared to Mind-Video, our model exhibits a smaller performance drop when the stimulus set is expanded to the ’Large’ scale, demonstrating its generalization capability. Notably, Wen’s results surpass those of Kupershmidt largely and are comparable to Mind-Video. This can be attributed to their approach of reconstructing a single frame from each video segment, simplifying the video reconstruction task into an image reconstruction task, thereby achieving superior performance on the retrieval metrics.
@@ 5.2 Ablation Study
Models Semantic-level ↑ Pixel-level ↑ ST-level
2-way-I 2-way-V VIFI-score SSIM PSNR Hue-pcc CLIP-pcc↑ EPE↓
w/o Semantic 0.679 0.766 0.523 0.097 8.005 0.737 0.123 8.719
w/o Structure 0.789 0.814 0.555 0.184 8.712 0.791 0.260 7.683
w/o Motion 0.674 0.789 0.585 0.136 8.611 0.715 0.376 6.374
Full Model 0.812 0.839 0.604 0.319 9.116 0.778 0.413 5.572
Table 6: Ablation study about our proposed decoders on subject 1 of CC2017 dataset. More results on subject 2 and 3 can be found in Appendix 12 and 13. 100 sample sets were constructed using the bootstrap method for hypothesis testing. Colors reflect statistical significance (Wilcoxon test for paired samples) compared to the Full Model. $p<0.0001$ (purple); $p<0.01$ (pink); $p<0.05$ (yellow); $p>0.05$ (green).
In this subsection, we conduct a detailed ablation study to assess the effectiveness of the three decoders we proposed and evaluate the impact of various hyperparameters on video reconstruction. First, we present the results obtained using the full model. Then, based on the full model, we eliminate the semantic decoder (w/o Semantic) and the structure decoder (w/o Structure) separately, replacing their outputs with random Gaussian noise. For the consistency motion generator, we replace it with 8 simple MLPs to model each frame individually (w/o Motion). Table 6 demonstrates that the removal of any decoder results in a significant decline in the model’s performance across nearly all metrics, which shows the efficacy of our proposed decoders. Notably, the Hue-pcc significantly increased after removing the structure decoder. We hypothesize that while fMRI contains structure and motion information, its low signal-to-noise ratio introduces noise that affects the generation quality of Sable Diffusion.
@@ 6.1 Have we truly decoded motion information from fMRI?
Following the work of Wang et al. (2022), we conduct shuffle tests on three subjects from the CC2017 dataset to evaluate whether the CMG accurately decodes motion information from fMRI, focusing on reconstructed videos with clear semantic decoding. Specifically, for each 8-frame reconstructed video clip from each subject, we shuffle the frame order 100 times randomly and compute spatiotemporal-level metrics on the original and shuffled frames. Subsequently, we estimate the P-value by the following formula: $P=\sum_{i=1}^{100}\delta_{i}/100$, where $\delta_{i}=1$ if the $i$-th shuffle outperforms the reconstruction result in the original order based on the metrics; otherwise, $\delta_{i}=0$. A lower P-value signifies a closer alignment between the sequential order of the reconstructed video and the ground truth. We repeat the shuffle test 5 times under conditions with and without the CMG, as illustrated in Figure 6.
Figure 6: The results of shuffle test on the CC2017 dataset. The experiment is repeated 5 times on 3 subjects, with the mean and std presented in subplots (a), (b), and (c), respectively. Paired t-tests with Bonferroni correction are performed, with significance denoted as $p<0.001(***)$, $p<0.01(**)$, $p<0.05(*)$, and $p>0.05(NS)$ for non-significant results.
It can be observed that the P-value of EPE is significantly lower than 0.05 when CMG is applied. However, although the P-value of CLIP-pcc is significantly smaller with CMG compared to without CMG, the P-value remains significantly greater than 0.05. To explain this, we further repeated the shuffle test on the reconstruction results’ noise ceiling (videos generated directly using the test set features). The results show that even for the noise ceiling, the P-value of CLIP-pcc remains significantly greater than 0.05. This indicates that: (1) we have indeed decoded motion information from fMRI, and (2) EPE is a more effective metric than CLIP-pcc for evaluating the model’s ability to decode motion information.
Figure 7: Ablation experiment results of fMRI guidance on the test sets of subject 1 and subject 2 from the CC2017 dataset.
To further validate whether the decoded motion information originates from fMRI guidance or the CMG’s autoregressive training, we removed the fMRI guidance during the CMG module’s training (w/o fMRI guidance) by replacing the cross-attention in the Spatial Module with self-attention, while keeping the rest of the architecture and hyperparameters unchanged. As shown in Figure 7, removing the fMRI guidance led to a significant deterioration in EPE, confirming that the proposed CMG effectively decodes motion information from fMRI. Additionally, comparing the removal of the entire CMG module (w/o Motion) with the removal of fMRI guidance (w/o fMRI guidance), we find that the latter accounts for the majority of the impact on EPE (i.e., 90% of the decrease in EPE can be attributed to the absence of fMRI guidance). This further underscores the critical role of fMRI guidance in accurately decoding motion information from brain signals.
@@ 6.2 Which brain regions are responsible for decoding different features, respectively?
To investigate voxels in which brain regions are responsible for decoding different features (semantic, structure, motion) during the fMRI-to-feature stage, we compute the voxel-wise importance maps in the visual cortex. Specifically, for a trained decoder, we multiply the weight matrix of the linear layers, then average the results across the feature dimension, and normalize them to estimate the importance weights for each voxel. A higher weight indicates that the voxel plays a more significant role in feature decoding. We project the importance maps of subject 1’s voxels from the CC2017 dataset onto the visual cortex, as depicted in Figure 8. To obtain ROI-wise importance maps, we calculate the average of the importance weights of voxels contained within each Region of Interest (ROI), with the results presented in Figure 9. The results from other subjects are presented in Appendix F.
@@
Figure 8 (a) indicates that high-level visual cortical areas (HVC, such as MT) contribute more significantly to the decoding of semantic feature, accounting for 60.5% of the total, as shown in Figure 9 (a). Figure 8 (c) and 9 (c) indicates that both LVC and HVC contribute to the decoding of motion information, with significant weight attributed to MT and TPOJ. This observation is consistent with previous work (Born & Bradley (2005)), which validates the function of MT and TPOJ in visual motion perception and processing.
We also identify the following findings in Figure 9: (1) MT shows significant activation for semantic decoding. This observation aligns with the functional segregation and interaction between the dorsal and ventral pathways during dynamic visual input processing (Ingle et al. (1982)). Specifically, the dorsal-dorsal pathway is associated with action control, whereas the ventral-dorsal pathway is involved in action understanding and recognition (Rizzolatti & Matelli (2003)). This finding aligns with the latter. (2) V1 is predominantly activated when decoding motion features, reflecting the visual system’s parallel processing capability. Motion information in the dorsal pathway does not strictly follow hierarchical processing (Zeki & Shipp (1988)). As noted by Nassi et al. (Nassi & Callaway (2009)), V1 directly projects motion-related information, such as direction and speed, to MT for further processing. For detailed neurobiological explanations, please refer to the Appendix G.